# Analytical Emulator for the Baryon EoM inside the Fuzzy Dark Matter Soliton from Machine Learning

### v1

@@ Abstract
An empirical baryon density profile can be included in the Schrödinger-Poisson (SP) equations to influence the fuzzy dark matter (FDM) soliton formation. However, to probe the effects of baryon on the other dynamical evolutions of the FDM soliton, its equation of motion (EoM) inside the corresponding FDM soliton is needed. In this paper, given an empirical baryon density profile, we first provide the cylinderical symmetric FDM soliton solution about the FDM density and the total potential of FDM and baryon. Then, we build an analytical baryon EoM from the obtained FDM density and total potential by machine learning . Finally, we check that this baryon EoM works as well as an empirical baryon density profile for the FDM soliton formation, with the fractional errors $\lesssim 0.04$. It should also work well for some other simple FDM soliton evolutions.
@@ I Introduction
In this paper, we propose a novel method to construct the baryon EoM inside the FDM soliton given a mock data of FDM and baryon density profile inside the FDM soliton. Usually, it’s very easy to reconstruct a function of one variable $f(x)$ according to data, such as reconstruction of equation of state of dark energy $w(z)$, Hubble parameter $H(z)$, luminosity distance $D_{L}(z)$ or cosmic reionization history $x_{\rm e}(z)$ by MCMC global fitting Collett:2019hrr , principal components analysis Huang:2017huh , Gaussian process Shafieloo:2012ht ; Seikel:2012uu ; Yang:2024rsy , artificial neural networks(ANNs) Wang:2019vxv or genetic algorithms (GAs) Arjona:2019fwb ; Nesseris:2012tt . As for the function of multiple variables f(x,y,)f(x,y,...), machine learning methods including GAs and ANNs are also the better choices for its reconstruction from data, such as the built of the analytical emulator for the linear matter power spectrum $P(k;\omega_{\rm b},\omega_{\rm m})$ by GAs Orjuela-Quintana:2022nnq ; Orjuela-Quintana:2023uqb ; Orjuela-Quintana:2024hha ; Bartlett:2023cyr and the gravitational wave surrogate models by ANNs Setyawati:2019xzw ; Khan:2020fso ; Thomas:2022rmc . Since the baryon EoM is supposed to be a function of FDM density and total potential $\rho_{\rm b,EoM}(\rho,\Phi)$, we will turn to GAs to build it too.
This paper is organized as follows. In Section II, we solve the SP system with an empirical baryon density profile background in cylindrical coordinates. In Section III, we extract a mock data of baryon density profile and then construct the analytical emulator for baryon EoM by GA. In Section IV, we check our baryon EoM by including it an enlarged SP system. Finally, a brief summary and discussion are provided in Section V.
@@ II FDM soliton solution with an empirical baryon density profile
where $m$ is the mass of the FDM particle, $\psi$ is its wavefunction, and $\Phi$ is the total gravitational potential, which is sourced by the FDM density $\rho=|\psi|^{2}$ and the baryon density $\rho_{\rm b}$. For the Milky Way, as Bar:2019bqz ; Launhardt:2002tx , we only consider the main contributions: a nuclear stellar disk (NSD) and a spherical nuclear cluster (NSC)
@@
where $r=\sqrt{R^{2}+z^{2}}$, $\bar{\rho}_{\rm NSD}=330M_{\odot}/{\rm pc}^{3}$, $\bar{\rho}_{\rm NSC}=3.3\times 10^{6}M_{\odot}/{\rm pc}^{3}$ and $n_{\rm NSC}=2$ for $r<6{\rm pc}$ and $\bar{\rho}_{\rm NSC}=9.0\times 10^{7}M_{\odot}/{\rm pc}^{3}$ and $n_{\rm NSC}=3$ for $r\geq 6{\rm pc}$.
@@
When $\lambda_{0}=\lambda_{1}$, it means that $\rho_{\rm b}$ and $M$ are equal to the baryon density and the FDM soliton mass of the Milky Way at the same time. Therefore, if $\lambda_{0}\neq\lambda_{1}$, we replace $\lambda_{0}$ with $\lambda_{1}$ and repeat the above three steps Tan:2024dne . We stop the iteration until $\lambda_{1}/\lambda_{0}-1>0.0001$ or $\lambda_{0}\approx\lambda_{1}\approx 6\times 10^{-7}$ for [^1] Here $m=10^{-22}~\rm{eV}/c^{2}$ just serve as an example for didactic purpose. Although different $m$ leads to different results in the following discussion, the method proposed in this paper could still hold. $m=10^{-22}~\rm{eV}/c^{2}$. In Fig. 1 and Fig. 2, we show the final solutions of $\rho_{\rm b}$ and $\Phi$, which are consistent with the results of  Bar:2019bqz . These solutions are dependent on $\tan\theta=\frac{z}{R}$, which will serve as mock data to construct the baryon EoM inside the FDM soliton.
Figure 1: FDM soliton density contours in the inner Milky Way, which is consistent with Fig. 2 of Bar:2019bqz . The corresponding dimensionless total baryon mass is $\tilde{M}_{\rm b}=1.03$.
@@
III Construction of baryon EoM by machine learning
@@
in this section, we look for a mathematical expression optimizing the goodness of fit to the given data of $\tilde{\rho}_{\rm b}$. First, we describe how the mock data of $\tilde{\rho}_{\rm b}$ for the fitting process was gathered. Then we construct the fitting function or analytical emulator for baryon EoM by GA.
@@ III.1 Mock data
Since Fig. 1 shows the distribution of $\rho(R,z)$, from it and by setting $\lambda\approx 6\times 10^{-7}$, we can extract $\tilde{\phi}^{2}(\tilde{R},\theta)$ for a given pair of $\{\tilde{R},\theta\leq 45^{\circ}\}$ or $\tilde{\phi}^{2}(\tilde{z},\theta)$ for a given pair of $\{\tilde{z},\theta>45^{\circ}\}$. Here, we consider that $\theta=0^{\circ},1^{\circ},...,90^{\circ}$ and $\Delta\tilde{R}=\Delta\tilde{z}=0.01$. However, we only sample $91\times 91$ $\tilde{\phi}^{2}$ points by $91\times 46$ pairs of $\{\tilde{R},\theta\leq 45^{\circ}\}$ and $91\times 45$ pairs of $\{\tilde{z},\theta>45^{\circ}\}$. Similarly, we also sample $91\times 91$ $\tilde{\Phi}$ points by the same pairs of $\{\tilde{R},\theta\leq 45^{\circ}\}$ and $\{\tilde{z},\theta>45^{\circ}\}$ from Fig. 2. Combing the corresponding $91\times 91$ $\tilde{\rho}_{\rm b}$ points retrieved from Eq. (II), our final dataset is a table of dimensions $8281\times 4$ with data points given as $\{\theta,\tilde{\phi}^{2},\tilde{\Phi},\tilde{\rho}_{\rm b}\}$. We quantify the goodness of fit of a given analytical expression for baryon EoM by the following function
$\%{\rm Acc}=\frac{100}{N}\sum_{i}^{N}\left|\frac{\lg[\tilde{\rho}_{{\rm b},i}]-\lg[\tilde{\rho}_{{\rm b},i,{\rm EoM}}]}{\lg[\tilde{\rho}_{{\rm b},i}]}\right|$ (8)
@@
III.2 Analytical emulator for baryon EoM
GA looks for a fitting formula to a dataset by evolving the population, each individual of which takes a form similar to Eq. (7). More precisely, in this paper, we set the analytical emulator for baryon EoM as
$\lg[\tilde{\rho}_{\rm b,EoM}(x,\theta)]$ $\equiv\lg 21551.2+f_{\rm EoM}(3.75395,\theta)-f_{\rm EoM}(x,\theta),$
$f_{\rm EoM}(x,\theta)$ $\equiv\sum_{i=1}^{7}a_{i}(\theta)g_{i}\{b_{i}(\theta)[x-c_{i}(\theta)]\}d_{i}(x),$
@@ III.2 Analytical emulator for baryon EoM
where $21551.2$ comes from the maximum value of the mock data of $\tilde{\rho}_{\rm b}$. The sum goes from 1 to 7 because the chromosome of each individual contains 7 genes and each gene is composed of several nucleotides $\{a_{i},b_{i},c_{i},d_{i},g_{i}\}$. $g_{i}$ is an operation in the grammar set $\{x,e^{x},-e^{-x}\}$ and, for simplicity, we fix them as
@@
$\{a_{i},b_{i},c_{i}\}$ are linear functions of $\sin\theta$ to characterize the cylindrical symmetry of $f_{\rm EoM}(x,\theta)$
@@
where $\{a^{\prime}_{i},b^{\prime}_{i},c^{\prime}_{i}\}\in[0,5]$ and $\{a^{\prime\prime}_{i},b^{\prime\prime}_{i},c^{\prime\prime}_{i}\}\in[-1,1]$ are random numbers. $d_{i}$ is a correction that can be fine-tuned to improve the goodness of fit
@@
We perform a stochastic search with GA code Arjona:2019fwb ; Orjuela-Quintana:2022nnq ; Orjuela-Quintana:2023uqb ; Orjuela-Quintana:2024hha and find a function with $\%{\rm Acc}=2.54844$, whose parameters are listed in Tab. 1. We also plot the accuracy of this analytical emulator for baryon EoM at different $\theta$ in Fig. 3. The accuracy is $\%4$, which means that our emulator can be used for some simple baryon evolutions inside the FDM soliton. In fact, the accuracy at $\theta\gtrsim 60^{\circ}$ can be further improved if the unphysical jump in mock data $\tilde{\rho}_{\rm b}$ due to the cutoff of Eq. (II) disappears. In Fig. 4, the comparison between $\tilde{\rho}_{\rm b,EoM}(x,\theta)$ (colored points) and $\tilde{\rho}_{\rm b}(x,\theta)$ (black points) at four angles $\theta=\{0^{\circ},30^{\circ},60^{\circ},90^{\circ}\}$ is shown respectively.
Table 1: Best-fit parameters for the baryon EoM emulator given in Eq. (III.2).
@@
Figure 3: Accuracy of the analytical emulator for baryon EoM as a function of $\theta$.
Figure 4: Comparison between the analytical emulator for baryon EoM $\tilde{\rho}_{\rm b,EoM}(x,\theta)$ (colored points) and the mock data $\tilde{\rho}_{\rm b}(x,\theta)$ (black points) at four angles $\theta=\{0^{\circ},30^{\circ},60^{\circ},90^{\circ}\}$ respectively. The unphysical jump in $\tilde{\rho}_{\rm b}(x,\theta)$ for larger $\theta$ comes from the cutoff of Eq. (II), which decreases the accuracy to some extent.
IV FDM soliton formation with an analytical baryon EoM
Note that the above analytical emulator for baryon EoM is constructed from dimensionless mock data $\{\theta,\tilde{\phi}^{2},\tilde{\Phi},\tilde{\rho}_{\rm b}\}$. Therefore, replacing the fixed baryon background $\tilde{\rho}_{\rm b}$ of Eq. (3) with an analytical baryon EoM $\tilde{\rho}_{\rm b,EoM}$ (Eq. (III.2)) leads to a new coupled system
$\begin{cases}\begin{aligned} &\tilde{\nabla}^{2}\tilde{\phi}=2(\tilde{\Phi}+\tilde{\gamma})\tilde{\phi},\\ &\tilde{\nabla}^{2}\tilde{\Phi}=\tilde{\phi}^{2}+\tilde{\rho}_{\rm b,EoM},\\ &\lg\tilde{\rho}_{\rm b,EoM}=4.3+f_{\rm EoM}(3.75,\theta)-f_{\rm EoM}(x,\theta),\end{aligned}\end{cases}$ (13)

### v2

# Analytical Emulator for the Baryon Density Distribution inside the Fuzzy Dark Matter Soliton from Machine Learning

@@ Abstract
An empirical baryon density profile can be included in the Schrödinger-Poisson (SP) equations to influence the fuzzy dark matter (FDM) soliton formation. However, to probe the effects of baryon on the other dynamical evolutions of the FDM soliton, its equation of motion (EoM) inside the corresponding FDM soliton is needed. In this paper, given an empirical baryon density profile, we first provide the cylindrical symmetric FDM soliton solution about the FDM density and the total potential of FDM and baryon. Then, instead of EoM, we build an analytical emulator (AE) for the baryon density distribution from the obtained FDM density and total potential by machine learning. Finally, we check that this AE works as well as an empirical baryon density profile for the FDM soliton formation, with the fractional errors $\lesssim 0.04$. It should also work as a kind of baryon EoM for some other simple FDM soliton evolutions.
@@ I Introduction
In this paper, instead of building the baryon EoM inside the FDM soliton directly, we propose a novel method to construct an analytical emulator (AE) for the baryon density distribution given a mock data of FDM and baryon density profile inside the FDM soliton. Usually, it’s very easy to reconstruct a function of one variable $f(x)$ according to data, such as reconstruction of equation of state of dark energy $w(z)$, Hubble parameter $H(z)$, luminosity distance $D_{L}(z)$ or cosmic reionization history $x_{\rm e}(z)$ by MCMC global fitting Collett:2019hrr , principal components analysis Huang:2017huh , Gaussian process Shafieloo:2012ht ; Seikel:2012uu ; Yang:2024rsy , artificial neural networks(ANNs) Wang:2019vxv or genetic algorithms (GAs) Arjona:2019fwb ; Nesseris:2012tt . As for the function of multiple variables f(x,y,)f(x,y,...), machine learning methods including GAs and ANNs are also the better choices for its reconstruction from data, such as the built of the analytical emulator for the linear matter power spectrum $P(k;\omega_{\rm b},\omega_{\rm m})$ by GAs Orjuela-Quintana:2022nnq ; Orjuela-Quintana:2023uqb ; Orjuela-Quintana:2024hha ; Bartlett:2023cyr and the gravitational wave surrogate models by ANNs Setyawati:2019xzw ; Khan:2020fso ; Thomas:2022rmc . Since AE for the baryon density distribution is supposed to be a function of FDM density and total potential $\rho_{\rm b,AE}(\rho,\Phi)$, we will turn to GAs to build it too.
This paper is organized as follows. In Section II, we solve the SP system with an empirical baryon density profile background in cylindrical coordinates. In Section III, we extract a mock data of baryon density profile and then construct AE for the baryon density distribution by GA. In Section IV, we check our baryon AE by including it in an enlarged SP system. Finally, a brief summary and discussion are provided in Section V.
@@ II FDM soliton solution with an empirical baryon density profile
where $m$ is the mass of the FDM particle, $\psi$ is its wavefunction, and $\Phi$ is the total gravitational potential, which is sourced by the FDM density $\rho=|\psi|^{2}$ and the baryon density $\rho_{\rm b}$. For the Milky Way, as Bar:2019bqz ; Launhardt:2002tx , we only consider the main contributions: a nuclear stellar disk (NSD) and a spherical nuclear stellar cluster (NSC)
@@
where $r=\sqrt{R^{2}+z^{2}}$, $\bar{\rho}_{\rm NSD}=330M_{\odot}/{\rm pc}^{3}$, $\bar{\rho}_{\rm NSC}=3.3\times 10^{6}M_{\odot}/{\rm pc}^{3}$ and $n_{\rm NSC}=2$ for $r<6{\rm pc}$, $\bar{\rho}_{\rm NSC}=9.0\times 10^{7}M_{\odot}/{\rm pc}^{3}$ and $n_{\rm NSC}=3$ for $r\geq 6{\rm pc}$ and NSC has an outer cut-off radius at $r=200{\rm pc}$.
@@
When $\lambda_{0}=\lambda_{1}$, it means that $\rho_{\rm b}$ and $M$ are equal to the baryon density and the FDM soliton mass of the Milky Way at the same time. Therefore, if $\lambda_{0}\neq\lambda_{1}$, we replace $\lambda_{0}$ with $\lambda_{1}$ and repeat the above three steps Tan:2024dne . We stop the iteration until $\lambda_{1}/\lambda_{0}-1>0.0001$ or $\lambda_{0}\approx\lambda_{1}\approx 6\times 10^{-7}$ for [^1] Here $m=10^{-22}~\rm{eV}/c^{2}$ just serve as an example for didactic purpose. Although different $m$ leads to different results in the following discussion, the method proposed in this paper could still hold. $m=10^{-22}~\rm{eV}/c^{2}$. In Fig. 1 and Fig. 2, we show the final solutions of $\rho$ and $\Phi$, which are consistent with the results of  Bar:2019bqz . These solutions are dependent on $\tan\theta=\frac{z}{R}$, which will serve as mock data to construct the baryon AE inside the FDM soliton.
Figure 1: FDM soliton density contours in the inner Milky Way, which is consistent with Fig. 2 of Bar:2019bqz . To compared with Fig. 2 of Bar:2019bqz easily, the density is also normalised to a reference value of $23.6M_{\odot}/{\rm pc}^{3}$ as Fig. 2 of Bar:2019bqz . The corresponding dimensionless total baryon mass is $\tilde{M}_{\rm b}=1.03$.
@@
III Construction of baryon AE by machine learning
@@
in this section, we look for a mathematical expression optimizing the goodness of fit to the given data of $\tilde{\rho}_{\rm b}$. First, we describe how the mock data of $\tilde{\rho}_{\rm b}$ for the fitting process was gathered. Then we construct the fitting function or AE for the baryon density distribution by GA.
@@ III.1 Mock data
Since Fig. 1 shows the distribution of $\rho(R,z)$, from it and by setting $\lambda\approx 6\times 10^{-7}$, we can extract $\tilde{\phi}^{2}(\tilde{R},\theta)$ for a given pair of $\{\tilde{R},\theta\leq 45^{\circ}\}$ or $\tilde{\phi}^{2}(\tilde{z},\theta)$ for a given pair of $\{\tilde{z},\theta>45^{\circ}\}$. Here, we consider that $\theta=0^{\circ},1^{\circ},...,90^{\circ}$ and $\Delta\tilde{R}=\Delta\tilde{z}=0.01$. However, we only sample $91\times 91$ $\tilde{\phi}^{2}$ points by $91\times 46$ pairs of $\{\tilde{R},\theta\leq 45^{\circ}\}$ and $91\times 45$ pairs of $\{\tilde{z},\theta>45^{\circ}\}$. Similarly, we also sample $91\times 91$ $\tilde{\Phi}$ points by the same pairs of $\{\tilde{R},\theta\leq 45^{\circ}\}$ and $\{\tilde{z},\theta>45^{\circ}\}$ from Fig. 2. Combing the corresponding $91\times 91$ $\tilde{\rho}_{\rm b}$ points retrieved from Eq. (II), our final dataset is a table of dimensions $8281\times 4$ with data points given as $\{\theta,\tilde{\phi}^{2},\tilde{\Phi},\tilde{\rho}_{\rm b}\}$. We quantify the goodness of fit of a given analytical expression for baryon AE by the following function
$\%{\rm Acc}=\frac{100}{N}\sum_{i}^{N}\left|\frac{\lg[\tilde{\rho}_{{\rm b},i}]-\lg[\tilde{\rho}_{{\rm b},i,{\rm AE}}]}{\lg[\tilde{\rho}_{{\rm b},i}]}\right|$ (8)
@@
III.2 AE for the baryon density distribution
In fact, construction of AE for the baryon density distribution belongs to a symbolic regression problem where both the model structure and parameters should be discovered from the data. In particular, the model structure can be fixed after a large number of generations by GA’s genetic operators, namely, crossover and mutation.
GA looks for a fitting formula to a dataset by evolving the population, each individual of which takes a form similar to Eq. (7). More precisely, in this paper, we set the analytical expression for baryon AE as
$\lg[\tilde{\rho}_{\rm b,AE}(x,\theta)]$ $\equiv\lg 21551.2+f_{\rm AE}(3.75395,\theta)-f_{\rm AE}(x,\theta),$
$f_{\rm AE}(x,\theta)$ $\equiv\sum_{i=1}^{7}a_{i}(\theta)g_{i}\{b_{i}(\theta)[x-c_{i}(\theta)]\}d_{i}(x),$
@@ III.2 AE for the baryon density distribution
where $21551.2$ comes from the maximum value of the mock data of $\tilde{\rho}_{\rm b}$, $3.75395$ is the value of $\tilde{\phi}^{2}-\tilde{\Phi}$ at $(\tilde{R}=0,\tilde{z}=0)$, functions $\{g_{i}(x)\}$ determine the model structure and $\{a_{i},b_{i},c_{i},d_{i}\}$ serve as the “amplitude”, “frequency”, “phase” and a correction respectively. The sum goes from 1 to 7 because the chromosome of each individual contains 7 genes and each gene is composed of several nucleotides $\{a_{i},b_{i},c_{i},d_{i},g_{i}\}$. $g_{i}$ is an operation in the grammar set $\{x,e^{x},-e^{-x}\}$ and, for simplicity, we fix them as
@@
$\{a_{i},b_{i},c_{i}\}$ are linear functions of $\sin\theta$ to characterize the cylindrical symmetry of $f_{\rm AE}(x,\theta)$
@@
where $\{a^{\prime}_{i},b^{\prime}_{i},c^{\prime}_{i}\}\in[0,5]$ and $\{a^{\prime\prime}_{i},b^{\prime\prime}_{i},c^{\prime\prime}_{i}\}\in[-1,1]$ are free parameters. $d_{i}$ is a correction that can be fine-tuned to improve the goodness of fit
@@
As shown in Fig. 4, each set of mock data (black points) includes three parts: the left part drops dramatically and behaves as $e^{-x}$, the middle part features a kind of plateau of slow growth and behave as $x$, the right part increases dramatically and behaves as $e^{x}$. Furthermore, the unphysical jump for larger $\theta$ divides the left part into two sub parts. When we describe each part (or sub part) with the combination of at least two corresponding elementary functions, there would be 88 $g_{i}(x)$. Since the combination of two linear functions can be replaced with another new linear function, we finally fixed 77 $g_{i}(x)$ as Eq. (III.2) before the code starts running. Due to the cylindrical symmetry, the “amplitude”, “frequency” and “phase” $\{a_{i},b_{i},c_{i}\}$ should depends on $\theta$. When we choose $\theta=0^{\circ}$ as the pivot angle, the dependence of $\{a_{i},b_{i},c_{i}\}$ on $\theta$ at the other angles can be simply included as Eq. (III.2). After many trials, we find that any combination of $g_{1}(x)$ and $g_{2}(x)$ can not fit the right part of mock data well. To improve the goodness of fit, we introduce corrections $\{d_{i}\}$ as Eq. (12) and fine tune their factors as $0.18$. In principle, GA can automatically discover the counterparts of Eq. (III.2), Eq. (III.2) and Eq. (12) during the stochastic search. However, in practice, we provide them in advance to reduce the computational cost.
We perform a stochastic search with GA code Arjona:2019fwb ; Orjuela-Quintana:2022nnq ; Orjuela-Quintana:2023uqb ; Orjuela-Quintana:2024hha and find a function with $\%{\rm Acc}=2.54844$, whose parameters are listed in Tab. 1. We also plot the accuracy of this baryon AE at different $\theta$ in Fig. 3. The accuracy is $4\%$, which means that our emulator can be used for some simple baryon evolutions inside the FDM soliton. In fact, the accuracy at $\theta\gtrsim 60^{\circ}$ can be further improved if the unphysical jump in mock data $\tilde{\rho}_{\rm b}$ due to the cutoff of Eq. (II) disappears. In Fig. 4, the comparison between $\tilde{\rho}_{\rm b,AE}(x,\theta)$ (colored points) and $\tilde{\rho}_{\rm b}(x,\theta)$ (black points) at four angles $\theta=\{0^{\circ},30^{\circ},60^{\circ},90^{\circ}\}$ is shown respectively.
Table 1: Best-fit parameters for the baryon AE given in Eq. (III.2).
@@
Figure 3: Accuracy of the baryon AE as a function of $\theta$.
Figure 4: Comparison between the baryon AE $\tilde{\rho}_{\rm b,AE}(x,\theta)$ (colored points) and the mock data $\tilde{\rho}_{\rm b}(x,\theta)$ (black points) at four angles $\theta=\{0^{\circ},30^{\circ},60^{\circ},90^{\circ}\}$ respectively. The unphysical jump in $\tilde{\rho}_{\rm b}(x,\theta)$ for larger $\theta$ comes from the cutoff of Eq. (II), which decreases the accuracy to some extent.
IV FDM soliton formation with the baryon AE
Note that the above baryon AE is constructed from dimensionless mock data $\{\theta,\tilde{\phi}^{2},\tilde{\Phi},\tilde{\rho}_{\rm b}\}$. Therefore, replacing the fixed baryon background $\tilde{\rho}_{\rm b}$ of Eq. (3) with an baryon AE $\tilde{\rho}_{\rm b,AE}$ (Eq. (III.2)) leads to a new coupled system
$\begin{cases}\begin{aligned} &\tilde{\nabla}^{2}\tilde{\phi}=2(\tilde{\Phi}+\tilde{\gamma})\tilde{\phi},\\ &\tilde{\nabla}^{2}\tilde{\Phi}=\tilde{\phi}^{2}+\tilde{\rho}_{\rm b,AE},\\ &\lg\tilde{\rho}_{\rm b,AE}=4.3+f_{\rm AE}(3.75,\theta)-f_{\rm AE}(x,\theta),\end{aligned}\end{cases}$ (13)